Question 1 (1 point) A train running along a straight track at 73 m/s is slowed uniformly to a stop in 10.6 s. Find the stopping distance.
Question 2 (1 point) In the first part of your trip, you drive 55 miles at an average speed of 53 miles/hour. In the second part of your trip, you drive 5 hours at an average speed of 67 miles/hour. What is average speed during your whole trip?
Question 3 (1 point) A box slides down an incline with uniform acceleration, It starts from rest and attains a speed of 39.4 m/s in 5.5 sec. What distance the box will move in 23 sec?
Question 4 (1 point) A stone is thrown straight downward with initial speed of 5.2 m/s from a height of 94 m. What is the time it takes to reach the ground? 10
Answer #1

1) v = u + at

0 = 73 + a*10.6

a = -6.88 m/s^2

v^2- u^2 = 2aS

0 - 73^2 = 2*-6.88*S

S = 387 m

so the distance moved is 387 m

2) average speed = total distance / time

= (55 + 67*5) / ( 55/53 + 5)

= 64.6 miles/hr

3) v = u + at

39.4 = 0 + a*5.5

a = 7.16 m/s^2

S = ut + 0.5at^2

S = 0 + 0.5*7.16*23^2 = 1894.8 m

4) H = ut + 0.5gt^2

94 = 5.2t + 0.5*9.8*t^2

4.9t^2 + 5.2t - 94 = 0

t = -5.2 + sqrt(5.2^2 + 4*4.9*94) / 9.8

= 3.88 s